# remarkRemark

### v1

@@ 7 Convergence of the bSQH scheme
As the sufficient decrease condition (27) holds for each $k\in{\mathbb{N}}$, the sequence $(J_{B}({\mathbf{u}}^{(k)}))_{k\in\mathbb{N}}$ monotonically decreasing and the boundedness from below by zero of $J_{B}$ implies convergence. Summing (27) over the first $m\in{\mathbb{N}}$ members of the corresponding sequences, we obtain $J_{B}({\mathbf{u}}^{(0)})-J_{B}({\mathbf{u}}^{(m)})>\sum_{k=0}^{m}\eta{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}^{(k+1)}-{\mathbf{u}}^{(k)}|\kern-1.07639pt|\kern-1.07639pt|}.$ By $J_{B}({\mathbf{u}}^{(0)})<\infty$ and the monotonic decrease of the sequence of cost function values, we have that $m\rightarrow\infty$ implies $\lim_{k\rightarrow\infty}{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}^{(k)}_{l}-{\mathbf{u}}^{(k+1)}_{l}|\kern-1.07639pt|\kern-1.07639pt|}=0.$ Thus, the sequence $({\mathbf{u}}^{(k)})_{k\in{\mathbb{N}}}$ admits a unique accumulation point $\bar{{\mathbf{u}}}\in U_{ad}$ and the lower semicontinuity of $J_{B}$ implies $J_{B}(\bar{{\mathbf{u}}})\geq\inf_{u\in U_{ad}}J_{B}({\mathbf{u}})$.
We remark that for both of the prior results only the Assumptions A1-A2 were required. Therefore, the bSQH scheme is also appropriate for an application to supervised learning problems, where the regularizer is discontinuous and given by (25). Further, opposed to the discrete PMP formulation, the convexity of sets $U_{l}$ is not a vital requirement for the applicability of Algorithm 1 either. As a result and in contrast to the classical backpropagation framework, our methodology could readily be used to train discrete-weight neural networks.
@@
We apply the definition of the generalized directional derivative (5) to (44) and use the sum rule of the generalized directional derivative [30] together with the Cauchy-Schwartz inequality to derive for all $v_{l}\in U_{l}$:
@@
00 0 $\leq\limsup_{k\rightarrow\infty}\left(\sqrt{e^{(k)}}+\epsilon_{k}\|w_{l}^{(k+1)}-u_{l}^{(k)}\|\right)$
@@
This together with the sum rule for the Clarke subdifferential [30] implies
@@
The conditions of Theorem 7.7 are easily satisfied in the full-batch setting due to Lemma 7.5 and when utilizing a reformulated network such as (23) together with an updating strategy corresponding to (24). While this theorem applies to neural networks regularized using Elastic-Net regularization, it also extends to other regularization approaches. Without considering a reformulated network, the maximization of the Hamiltonian function becomes more complex as the activation occurs after the affine transformation. Nevertheless, the property (40) can be enforced by taking into account that for $\epsilon_{k}$ sufficiently large, the augmented Hamiltonian becomes strictly concave [40]. Thus, methods such as quasi-Newton methods can determine a close approximation of the true global maximum of the augmented Hamiltonian within a few iterations and thereby enable convergence of the full-batch bSQH in these cases as well. In the mini-batch framework and due to the results of Lemma 5.1, the property (40) can be enforced at least asymptotically by successively increasing the mini-batch size over the course of training. We will validate numerically that the property (40) can similarly be enforced for a fixed mini-batch size by combining it with the approach (bSQHω), where $\zeta=1$.
@@ 8 Numerical experiments
where $m_{l}$ is the dimension of the underlying parameter space $U_{l}$. This investigation is split into three parts. First, the full-batch variant of Algorithm 1 is examined with both the sparsity enforcing Elastic-Net regularization and the non-continuous $L^{0}$-’norm’ based regularizier (25). In both cases, we also investigate the impact of the selection strategies (SQHζ) and (bSQHω) on the computational overhead of the bSQH scheme. The second part of our numerical investigation is concerned with the mini-batch variant of Algorithm 1. In particular, we investigate the impact of the mini-batch approach on the stability of the algorithm’s convergence and demonstrate the computational advantages of the mini-batch approach compared to its full-batch counterpart. In the third and last part of this section, we present an application of the bSQH method in its mini-batch setting to a multi-classification task for medical imaging. These experiments substantiate our claim that the results of this work still hold in the case of network architectures using ReLU activation functions and max-pooling operations, where the differentiability assumptions in A2 are not satisfied. In all of the upcoming experiments, the weights of the network architectures are initialized using Xavier-normal initialization [41], while at the start of training the bias vectors are set to zero. An appropriate parameter update $w\in U_{l}$ for every layer at each iteration of the bSQH scheme is generated corresponding to the updating strategy (24) which readily applies in the framework of a LeNet-5 architecture and guarantees that the property (40) is satisfied exactly in the full-batch framework. This reformulation-based updating approach is different to the L-BFGS-based [42] updating technique commonly used in the PMP-based machine learning framework; e.g. see [12, 13] and enables training neural networks with non-continuous regularizers such as the $L^{0}$-’norm’, which directly penalizes the number of non-zero parameters of a network.
@@ 8.1 Sparse network training
For our investigation in the full-batch setting, we train a standard LeNet-5 architecture to classify handwritten digits in the MNIST dataset [43]. The network structure is chosen as proposed in [2] and uses tanh activations together with average pooling layers. In our experiments, we consider as loss function the cross-entropy loss and test our algorithm for both of the sparsity enforcing regularizers, Elastic-Net and (25), where we set $\alpha=0.8$ and $\rho=1e-04$. The MNIST dataset consists of $60.000$ images for testing and $10.000$ images for training, where for our numerical investigation, we extract a suitable subset with $|B|=20.000$ from the dataset dedicated to testing and apply a data normalization as a preprocessing step. The LeNet-5 is then trained in a full-batch fashion using a maximum of $k_{max}=1500$ iterations of Algorithm 1, where we set $\eta=10e-10$ and initialize the augmentation parameter with $\epsilon_{0}=1$. To investigate the capability of the parameter selection strategy (bSQHω), we compare it with the selection technique (SQHζ), which was also used in [13]. The corresponding training results averaged over multiple different weight initializations are reported in Table 1 and Table 2, together with the corresponding standard deviation. Here ’ACC’ refers to the test accuracy based on $10.000$ images in the test set and ’Linesearch steps’ refers to the average total number of inference steps necessary to determine a suitable augmentation parameter $\epsilon_{k}$ in each iteration. Further, we report the average percentage level of sparsity in the parameters of the trained models in Table 1 and Table 2. For the parameter setting of the full-batch scheme, we choose $\zeta=0.01$, $\mu=7$, similar to [13], which can be interpreted as a coarse grid search for an appropriate augmentation parameter. On the one hand, this choice allows large updates in the learning parameters, positively affecting the algorithm’s convergence on an iteration basis. On the other hand, it can result in an underestimation of $\hat{\epsilon}_{k}$ and thus a higher number of optimization steps discarded in the line search procedure. One can see in Table 1 and Table 2 that the latter effect is mitigated by the proposed strategy (bSQHω). In particular, our proposed estimation strategy for $\epsilon_{k}$ more than halves the required number of optimization steps compared to the standard SQH (SQHζ) approach by simultaneously retaining the testing accuracy and loss level achieved with (SQHζ). Further, since Theorem 7.3 together with Lemma 7.5 guarantees that the bSQH scheme imposes a decrease on the Loss function method even in the non-continuous setting of the sparsity enforcing regularizer (25) it can be readily applied for generating sparse CNNs. In the given case, the application of the bSQH scheme resulted in a LeNet-5 which uses only about $15\%$ of its parameters for classification by simultaneously achieving a level of Accuracy on the test set, which is similar to that of a network using about $60\%$ of its parameters.
@@ 8.2 Mini-batch bSQH training
In the following, the mini-batch variant of Algorithm 1 is investigated. For this purpose the previously defined Elastic-Net regularized problem is considered and the standard LeNet-5 is trained for a maximum of $k_{max}=2000$ iterations and different mini-batch sizes, that is, $|B_{k}|$ $=512$,  10241024\;10241024,  20482048\;20482048,  20.00020.000\;20.00020.000, where the latter corresponds to the full-batch mode of Algorithm 1. For the hyperparameters of the bSQH scheme, we set $\mu=1.1$ and determine estimates for the augmentation parameter at each iteration using (bSQHω), where $\omega=5$ and $\zeta=1$. This choice is motivated by the observation, that a reduction of the variance in $\epsilon_{k}$ leads to a reduced variance in the mini-batch augmented HP functions which carries over to the updating step of the network parameters. Further, the resulting damped variation of the augmentation parameter prevents also a constant decrease of the augmentation weight over the course of training, which would otherwise lead to overfitting to the individual mini-batches at each iteration. The learning curves are compared based on the relative GPU time required for training, that is, $T_{\%}$ denotes the percentage of training time required relative to the most time-consuming training process, which in the given case is the full-batch setting. In Figure 2 and for small mini-batch sizes, the variance induced by the mini-batch approach is apparent. However, when increasing the mini-batch size, due to Lemma 5.1 and the law of large numbers, the noise in the full-batch loss and validation accuracy is reduced, resulting in an almost monotonically decreasing sequence of cost functional values, similar to the full-batch setting. This provides evidence that the decrease properties of the full-batch bSQH algorithm corresponding to Lemma 7.5 hold at least approximately in the mini-batch setting. Based on our results, the mini-batch approach of bSQH has favourable convergence properties with respect to the GPU time and thus yields an appropriate learning strategy for training CNNs in a stochastic framework.
@@
$\Delta h:=\sum_{l=0}^{L-1}\max_{w\in U_{l}}\;{\cal H}^{B}_{l,\epsilon_{k}}({\mathbf{y}}_{l}^{(k)},w,u^{(k)}_{l})-{\cal H}^{B}_{l,\epsilon_{k}}({\mathbf{y}}_{l}^{(k)},u^{(k+1)}_{l},u^{(k)}_{l}),\qquad\Delta{\mathbf{u}}:=\sum_{i=k-5}^{k}{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}^{(i+1)}-{\mathbf{u}}^{(i)}|\kern-1.07639pt|\kern-1.07639pt|},$
with $k=2000$. While the value of $\Delta h$ serves as an estimate for (40), we have that $\Delta{\mathbf{u}}$ enables investigating the convergence properties of the parameter sequence $({\mathbf{u}}^{(k)})$ generated by the bSQH scheme. For our mini-batch framework with fixed mini-batch sizes, we observe in Table 3 that the conditions for Theorem 7.7 are satisfied at least approximately. Further, contributed to Lemma 5.1 we observe that these convergence properties are improved if the mini-batch size is increased.
@@ 8.3 The bSQH method for classification of medical images
In this Section, we apply the mini-batch version of the bSQH method to the OrganAMNIST dataset, which is provided as part of the medMNIST database [44, 45] and consists of a collection of computed tomography (CT) images of organs; see Figure 3 for samples with corresponding labels. The dataset consists of 58,850 images which are split into eleven categories. For solving the multiclassification task, the dataset is split into $|B|=34.581$ samples for training, $6.491$ images for validation and $17.778$ for testing.
@@
The CNN architecture we consider is a modification of the LeNet-5 architecture; see Table 4. These modifications involve an increase in the number of channels for each of the two convolutional layers, to allow for the extraction of more complex features, and the addition of trainable batch normalization operation (BN), aiding to reduce the internal covariate shift between layers. Each average-pooling operation is exchanged with a max-pooling operation to capture local feature information better. Additionally, we exchange the tanh activation with the widely used ReLU function which is much simpler computationally and introduces a sparsity into the feature maps by zeroing out negative activations, leading to faster forward and backward passes through the network and reduced memory requirement.
@@
[…]
@@ Conclusion
Convergence results for the bSQH scheme were presented and the effectiveness of its mini-batch variant was investigated numerically. Further, it was shown that the proposed bSQH algorithm in its full-batch and mini-batch mode had superior computational performance compared to a similar existing PMP-based scheme. Results of numerical experiments demonstrated the ability of the bSQH algorithm for sparse training of CNNs in image classification tasks by including a sparsity enforcing $L^{0}$-regularization.

### v2

@@ 7 Convergence of the bSQH scheme
[…]
We remark that for both of the prior results, only the Assumptions A1-A2 and the boundedness from below of $J_{B}$ were required. Therefore, the bSQH scheme is also appropriate for an application to supervised learning problems, where the regularizer is discontinuous and given by (24). Further, opposed to the discrete PMP formulation, the convexity of sets $U_{l}$ is not a vital requirement for the applicability of Algorithm 1 either. As a result, and in contrast to the classical backpropagation framework, our methodology could readily be used to train discrete-weight neural networks.
@@
We apply the definition of the generalized directional derivative (5) to (43) and use the sum rule of the generalized directional derivative [11] together with the Cauchy-Schwarz inequality to derive for all $v_{l}\in U_{l}$:
@@
$0$ $\leq\limsup_{k\rightarrow\infty}\left(\sqrt{e^{(k)}}+\epsilon_{k}\|w_{l}^{(k+1)}-u_{l}^{(k)}\|\right)$
@@
This, together with the sum rule for the Clarke subdifferential [11], implies
@@
The conditions of Theorem 7.7 are easily satisfied in the full-batch setting due to Lemma 7.5 and when using a reformulated network such as (22) together with an updating strategy corresponding to (23). While this theorem applies to supervised learning problems with Elastic-Net regularization, it also extends to other regularization approaches. Without considering a reformulated network, the maximization of the Hamiltonian function becomes more complex as the activation occurs after the affine transformation. Nevertheless, the property (39) can be enforced by taking into account that for $\epsilon_{k}$ sufficiently large, the augmented Hamiltonian becomes strictly concave [40]. Thus, methods such as quasi-Newton methods can determine a close approximation of the true global maximum of the augmented Hamiltonian within a few iterations and thereby enable convergence of the full-batch bSQH in these cases as well. In the mini-batch framework and due to the results of Lemma 5.1, the property (39) can be enforced at least asymptotically by successively increasing the mini-batch size throughout training. We will validate numerically that the property (39) can similarly be enforced for a fixed mini-batch size by combining it with the approach (bSQHω), where $\zeta=1$.
@@ 8 Numerical experiments
where $m_{l}$ is the dimension of the underlying parameter space $U_{l}$. This investigation is split into three parts. First, the full-batch variant of Algorithm 1 is examined with both the sparsity-enforcing Elastic-Net regularization and the non-continuous $L^{0}$-’norm’ based regularizier (24). In both cases, we also investigate the impact of the selection strategies (SQHζ) and (bSQHω) on the computational overhead of the bSQH scheme. The second part of our numerical investigation is concerned with the mini-batch variant of Algorithm 1. In particular, we investigate the impact of the mini-batch approach on the stability of the algorithm’s convergence and demonstrate the computational advantages of the mini-batch approach compared to its full-batch counterpart. In the third and last part of this section, we present an application of the bSQH method in its mini-batch setting to a multi-classification task for medical imaging. These experiments substantiate our claim that the results of this work still hold in the case of network architectures using ReLU activation functions and max-pooling operations, where the differentiability assumptions in A2 do not apply. In all of the upcoming experiments, the weights of the network architectures are initialized using Xavier-normal initialization [18], while at the start of training, the bias vectors are set to zero. An appropriate parameter update $w\in U_{l}$ for every layer at each iteration of the bSQH scheme is generated corresponding to the updating strategy (23) which readily applies in the framework of a LeNet-5 architecture and guarantees that the property (39) is satisfied exactly in the full-batch framework. This reformulation-based updating approach is different to the L-BFGS-based [30] updating technique commonly used in the PMP-based machine learning framework; e.g. see [28, 22] and enables training neural networks with non-continuous regularizers such as the $L^{0}$-’norm’, which directly penalizes the number of non-zero parameters of a network.
@@ 8.1 Sparse network training
For our investigation in the full-batch setting, we train a standard LeNet-5 architecture to classify handwritten digits in the MNIST dataset [13]. The network structure is chosen as proposed in [27] and uses $\tanh$ activations together with average pooling layers. In our experiments, we consider cross-entropy loss and test our algorithm for both of the sparsity-enforcing regularizers, Elastic-Net and (24), where $\alpha=0.8$ and $\rho=1e-04$.
@@
The MNIST dataset consists of $60.000$ images for testing and $10.000$ images for training. For our numerical investigation, we extract a suitable subset of $20.000$ images from the testing dataset and apply data normalization as a preprocessing step. The LeNet-5 is then trained in a full-batch fashion using a maximum of $k_{max}=1500$ iterations of Algorithm 1, where we set $\eta=10e-10$ and initialize the augmentation parameter with $\epsilon_{0}=1$. To investigate the capability of the parameter selection strategy (bSQHω), we compare it with the selection technique (SQHζ), which was also used in [22]. The corresponding training results averaged over multiple, differently seeded weight initializations are reported in Table 1 and Table 2, together with the corresponding standard deviation. Here, ’ACC’ refers to the test accuracy based on $10.000$ images in the test set and ’Linesearch steps’ refers to the average total number of inference steps necessary to determine a suitable augmentation parameter $\epsilon_{k}$ in each iteration. Further, we report the average percentage level of sparsity in the parameters of the trained models in Table 1 and Table 2. For the parameter setting of the full-batch scheme, we choose $\zeta=0.01$, $\mu=7$, similar to [22], which can be interpreted as a coarse grid search for an appropriate augmentation parameter. On the one hand, this choice allows large updates in the learning parameters, positively affecting the algorithm’s convergence on an iteration basis. On the other hand, it can result in an underestimation of $\hat{\epsilon}_{k}$ and thus a higher number of optimization steps discarded in the line search procedure. One can see in Table 1 and Table 2 that the latter effect is mitigated by the proposed strategy (bSQHω). In particular, our proposed estimation strategy for $\epsilon_{k}$ more than halves the required number of optimization steps compared to the standard SQH (SQHζ) approach by simultaneously retaining the testing accuracy and loss level achieved with (SQHζ). Further, since Theorem 7.3 together with Lemma 7.5 guarantees that the bSQH scheme imposes a decrease on the loss function method even in the non-continuous setting of the sparsity enforcing regularizer (24) it can be readily applied for generating sparse CNNs. In the given case, the application of the bSQH scheme resulted in a LeNet-5 which uses only about $15\%$ of its parameters for classification by simultaneously achieving a level of Accuracy on the test set, which is similar to that of a network using about $60\%$ of its parameters.
@@ 8.2 Mini-batch bSQH training
In the following, the mini-batch variant of Algorithm 1 is investigated. For this purpose the previously defined Elastic-Net regularized problem is considered and the standard LeNet-5 is trained for a maximum of $k_{max}=2000$ iterations and different mini-batch sizes, that is, $|B_{k}|$ =512=512,  1024\;1024,  2048\;2048,  20.000\;20.000, where the latter corresponds to the full-batch mode of Algorithm 1. For the hyperparameters of the bSQH scheme, we set $\mu=1.1$ and determine estimates for the augmentation parameter at each iteration using (bSQHω), where $\omega=5$ and $\zeta=1$. This choice is motivated by the observation that a reduction of the variance in $\epsilon_{k}$ leads to a reduced variance in the mini-batch augmented HP functions, which carries over to the updating step of the network parameters. Further, the resulting damped variation of the augmentation parameter also prevents a constant decrease of the augmentation weight over the course of training, which would otherwise lead to overfitting to the individual mini-batches at each iteration. The learning curves are compared based on the relative GPU time required for training, that is, $T_{\%}$ denotes the percentage of training time required relative to the most time-consuming training process, which in the given case is the full-batch setting. In Figure 2 and for small mini-batch sizes, the variance induced by the mini-batch approach is apparent. However, when increasing the mini-batch size, due to Lemma 5.1 and the law of large numbers, the noise in the full-batch loss and validation accuracy is reduced, resulting in an almost monotonically decreasing sequence of cost functional values, similar to the full-batch setting. This provides evidence that the minimizing properties of the full-batch bSQH algorithm corresponding to Lemma 7.5 hold at least approximately in the mini-batch setting. Based on our results, the mini-batch approach of bSQH has favourable convergence properties with respect to the GPU time and thus yields an appropriate learning strategy for training CNNs in a stochastic framework.
@@
[…]
with $k=2000$. While the value of $\Delta h$ serves as an estimate for (39), we have that $\Delta{\mathbf{u}}$ enables investigating the convergence properties of the parameter sequence $({\mathbf{u}}^{(k)})$ generated by the bSQH scheme. For our mini-batch framework with fixed mini-batch sizes, we observe in Table 3 that the conditions for Theorem 7.7 are satisfied at least approximately. Further, contributed to Lemma 5.1, we observe that these convergence properties are improved if the mini-batch size is increased.
@@ 8.3 The bSQH method for classification of medical images
In this Section, we apply the mini-batch version of the bSQH method to the OrganAMNIST dataset, which is provided as part of the medMNIST database [42, 43] and consists of a collection of computed tomography (CT) images of organs; see Figure 3 for samples with corresponding labels. The dataset consists of $58.850$ images, which are split into eleven categories. For solving the multiclassification task, the dataset is split into $|B|=34.581$ samples for training, $6.491$ images for validation and $17.778$ for testing.
@@
The CNN architecture we consider is a modification of the LeNet-5 architecture; see Table 4. These modifications involve an increase in the number of channels for each of the two convolutional layers, to allow for the extraction of more complex features, and the addition of trainable batch normalization operations (BN), aiding to reduce the internal covariate shift between layers. Each average-pooling operation is exchanged with a max-pooling operation to capture local feature information better. Additionally, we exchange the $\tanh$ activation with the widely used ReLU function, which is much simpler computationally and introduces sparsity into the feature maps by zeroing out negative activations, leading to faster forward and backward passes through the network and reduced memory requirement.
@@
The empirical risk minimization problem is formulated by combining a weighted cross entropy loss function with the sparsity enforcing regularizer (24) where $\alpha=0.99$ and a regularization weight $\rho=7.5\mathrm{e}{-03}$ is chosen. The weighting of the loss function is chosen to compensate for a moderate data imbalance within the training set. That is for the class $i$ we choose a weight as follows $w_{i}=\exp\left(\frac{|B|}{11\cdot c_{i}}\right)$, where $c_{i}$ denotes the total number of samples corresponding to the $i$-th class. The mini-batch bSQH method is applied with a maximum number of $k_{\max}=8.000$ iterations, using $\mu=1.1$ and the strategy (bSQHω) with $\omega=7,\zeta=1$ together with different mini-batch sizes $|B_{k}|=64,\;128,\;256$. We report in Table 5 the maximum testing accuracy (ACC) achieved throughout training, together with the corresponding values of the area under the receiver operating characteristic curve (AUC). Each result is the average over multiple weight initialization and is always reported together with the corresponding standard deviations. The AUC, in contrast to the ACC, is a threshold-free metric and thus less sensitive to class imbalance. Its values range from $0.5$, representing a random classifier, to $1.0$, which indicates a perfect classifier. In addition to the above evaluation metrics, we report the average percentage sparsity of the trained parameters corresponding to the reported ACC and AUC values. The confusion matrix for the best classifier results corresponding to Table 5 is given in Figure 4. The confusion matrix reveals that, in particular, the classes corresponding to CT images of the kidney cause a decrease in the network’s accuracy. This is due to strong similarities in the images of ’kidney-left’ and ’kidney-right’, possibly calling for a more elaborate network architecture. Nevertheless, and in particular, for a LeNet-5 type network architecture, the achieved results are satisfactory and can keep up with the benchmark results reported in [43]. In this context, it is important to emphasize that these results come from a model with a high level of sparsity that has only about $48.000$ trainable parameters. In contrast, models such as ResNet-18 have over a million trainable parameters and only achieve a $3\%$ increase in accuracy on the given dataset.
@@ Conclusion
Convergence results for the bSQH scheme were presented, and the effectiveness of its mini-batch variant was investigated numerically. Further, it was shown that the proposed bSQH algorithm in its full-batch and mini-batch modes has superior computational performance compared to a similar existing PMP-based scheme. Results of numerical experiments demonstrated the ability of the bSQH algorithm for sparse training of CNNs in image classification tasks by including a sparsity-enforcing $L^{0}$-regularization.